OpenStudy (anonymous):

Running 4/3 of his usual speed, A person improves his timing by 10 minutes. Find his usual time to cover the distance.

OpenStudy (amistre64):

s = d/t

OpenStudy (amistre64):

(4/3)Ns = d/(Nt-10)

OpenStudy (amistre64):

Ns = d/(Nt)

OpenStudy (amistre64):

(4/3)Ns = d/(Nt-10) Ns = d/(Nt) (4/3)d/(Nt) = d/(Nt-10)

OpenStudy (amistre64):

cross multiply (4/3)d(Nt-10) = d(Nt) d(Nt(4) - 40) ----------- = d(Nt) ; /d and *3 3 Nt(4) - 40 = Nt(3) Nt(4-3) = 40

OpenStudy (amistre64):

his Normalt time is 40 minutes

OpenStudy (amistre64):

..... maybe lol

OpenStudy (anonymous):

yes....... since time is inversely propotional to speed let his earlier time be x min so current time=3x/4 min =>x- (3x/4)= 10 =>x=40